determine the slope and coordinates of the y-intercept. no need to graph
-4x-3y=-6

Answer by checkley77(12844)

-4x-3y=-6
-3y=4x-6
y=-4x-3-6/-3
y=4x/3+2
slope=4/3
coordinates(0,2)